A 580-ft wide river spanned by a bridge, which has 1/6 of its length on the east bank and 1/7 of its length on the west bank. How long was the bridge?
Answer by nerdybill(7384)

A 580-ft wide river spanned by a bridge, which has 1/6 of its length on the east bank and 1/7 of its length on the west bank. How long was the bridge?
.
Let x = length of bridge
then
580 = x - (1/7 + 1/6)x
580 = x - (6/42 + 7/42)x
580 = x - (13/42)x
580 = x(1 - (13/42))
580 = x(42/42 - 13/42)
580 = x(29/42)
580(42/29) = x
20(42) = x
840 feet = x